a body moves in south for 10 s at the speed of 10 m/s.it then start moving east at the speed of 20 m/s for 10 s.the average velocity is?the average speed is?the total displacement is?

Grade:11

1 Answers

Aawadhut prakashrao Tekale
5 years ago
    body is moved in south 10m/s for 10s  
  •      so 1s =10m........(1)
  • 10s=10×10=100m
  • then the body move toward east for20m/s for 10s
  • so from(1)like
  • 10s=10s×20m=200m
  • then by hypotenous we can find displacement
  • di=223.60
average speed=total distance/time=100m+200m=300
  • and time=10s+10s=20s
  • now total dis/time=300/20
  • =15m/s
  • ave.velocity=Total dis/time=
  • 223.60/20
  • =11.18m/s 
  • This are the answer